pooja195 (pooja195):

How would I graph a line with a slope of 2/3?

OpenStudy (danjs):

start at the origin as your first point

OpenStudy (danjs):

slope = (change in y value) / change in x value) =2/3

OpenStudy (nincompoop):

slope = change in y divided by change in x m = 2/3

OpenStudy (nincompoop):



OpenStudy (nincompoop):

You don't have to start from (0,0). You can start anywhere, but it is just convenient if we did it from (0,0)

OpenStudy (danjs):

just use y = (2/3)*x and pick values for x to get points(x,y)

OpenStudy (nincompoop):

\(\sf slope = \frac{\Delta y}{\Delta x} = \frac{y_2-y_1}{x_2-x_1} = \frac{2-0}{3-0}=\frac{2}{3} \)

pooja195 (pooja195):

Thank you @nincompoop and @DanJS it seems alot more easier :)

OpenStudy (danjs):

If you want to move the line up or down then use any number b in, y = (2/3)*x + b

OpenStudy (danjs):

welcome

OpenStudy (nincompoop):

@DanJS don't get too ahead with the point-slope form, she is merely learning slope itself

OpenStudy (danjs):

oh, beg my pard

pooja195 (pooja195):

actually I know all this stuff I just dont know how to graph rise over run which now get :)

OpenStudy (nincompoop):

>.< rise/run is just a mnemonic devise

OpenStudy (nincompoop):

I'd be damned if you knew linear equation already and not know how graph a slope